# Multilingual Mechanism Analysis and Usage with MoE: Scaling Up to 50 Languages in medicine

### v1

@@ 5 Related Work
Mixture-of-Experts and Sparse Upcycling  Mixture-of-Experts (MoE) models differ from traditional dense models by activating only a subset of parameters, or ”experts,” for each input token. This selective activation reduces computational costs while preserving model capacity. First introduced by Shazeer et al. (2017), MoEs have since evolved in models like GShard (Lepikhin et al., 2020), Switch Transformers (Fedus et al., 2022), and Mixtral (Jiang et al., 2024), all of which have shown enhanced performance in large-scale language tasks. Recent advancements, such as Sparse Upcycling (Komatsuzaki et al., 2022), have efficiently initialized MoEs by replicating pre-trained dense model parameters across MoE layers, specifically duplicating feed-forward network (FFN) parameters for each expert. BTX (Sukhbaatar et al., 2024) extends this by upcycling multiple specialized models, further refining MoE efficiency and specialization.
Multilingual Medical LLMs  Recent research has increasingly focused on expanding multilingual medical data and developing models tailored to linguistically diverse populations. A LLM integrated with machine translation has been proposed to address the scarcity of multilingual medical datasets (Gangavarapu, 2024). Furthermore, multilingual medical corpora and benchmarks across six languages have been developed (Qiu et al., 2024; Wang et al., 2024). Our work builds on these efforts by investigating multilingual patterns and expanding it to 50 languages.
Multilingual Capability Enhancement Research on Multilingual Models has primarily aimed to enhance multilingual capabilities and understand their mechanisms. Efforts have improved performance through translation (Liang et al., 2023; Huang et al., 2023) and cross-lingual alignment (Salesky et al., 2023). Techniques like cross-lingual transfer (Kim et al., 2017; Lin et al., 2019) and continuous training in specific languages have further advanced LLMs (Cui et al., 2023), while training from scratch shows potential (Muennighoff et al., 2023). Despite these advancements, many methods are limited by task specificity or extensive data requirements, leaving the multilingual mechanisms of LLMs insufficiently explored. Recent models (Zhao et al., 2023; Nguyen et al., 2024) exhibit strong multilingual abilities without explicit language alignment. Tang et al. (2024) and Zhao et al. (2024) employ neuron analysis method (Mu & Andreas, 2020) to investigate the mechanism, but they typically focus on fewer than seven languages. Additionally, the complexities of neuron disassembly further limit the applicability of these findings.
@@ 6 Conclusion
In conclusion, this work advances the development of multilingual medical LLMs by addressing key challenges in data scarcity and model scalability. We first construct a high-quality medical dataset covering 12 major languages. Subsequently, we propose Hybrid-$k$ routing to explore multilingual training in a modular manner in MoE. Based on the Hybrid-$k$ routing, we propose a circuits-based paradigm for interpreting information flow in multilingual contexts. The circuit analysis reveale the ”Spread Out in the End” mechanism, based on which we introduced the Post-MoE architecture and demonstrate its superior multilingual capabilities. Furthermore, by introducing language family experts, we efficiently extend Post-MoE’s capabilities to 50 languages with limited data, demonstrating scalability without additional parameters. It is worth noting that the proposed method can be easily extended to support a broader range of languages, which we leave as straightforward future work. These contributions lay the groundwork for broader multilingual expansion in healthcare applications, promoting more equitable access to expert medical knowledge globally.
@@
Future Work  The Language-Specific experts and Hybrid routing are not limited to multilingual scenarios and provides a modular approach for exploring the internal mechanisms of MoE. By aligning multiple related tasks with a single expert and using experts as modules for model training, it reduces both parameter count and computational load while maintaining good performance. The Mixture of Language Family Experts in the Post-MoE architecture offers an approach for applying the model to a variety of tasks, enabling efficient expansion to more task scenarios.
@@ A.2 Dataset Procession
Adhering to the established data processing standards, we segmented sentences into chunks, filtered them for medical relevance, and reformatted them into a QA format with prompt shown in App. C.1. The processing details are as follows:
@@
A.3 Benchmark Across 12 Major Languages
@@ A.4 More Experimental Analysis on the Dataset
For Baselines, We select Gemma2-2B (Riviere et al., 2024), Phi-3-mini-4k (Abdin et al., 2024) and Qwen2-7B (Yang et al., 2024) as base models. We simultaneously selected the closed-source models, Claude3 Opus (Anthropic, 2024), and GPT-4o (Achiam et al., 2023), as well as highly competitive open-source models MMed-Llama-3-8B (Qiu et al., 2024), LLama3-OpenBioLLM-8B (Ankit Pal, 2024), and LLama3-8B (Dubey et al., 2024) as baselines. For Training, We use full-parameter fine-tuning with a learning rate of 1e-5 for the 7B and 9B models, and a learning rate of 1e-4 for smaller models. All model use a batch size of 32 and a sequence length of 4096 for training on 8x NVIDIA A800-SXM4-80GB GPUs.
@@
As shown in Tab. 8, the performance of each model significantly improved after training with 12 major languages data, further validating the effectiveness of data quality. The fine-tuned models excelled in the medical domain, significantly surpassing open-source medical models with similar parameter counts. Additionally, the medical model fully fine-tuned based on Gemma2-9B performed nearly as well as large-scale closed-source models.
A.5 Minor Languages Dataset&Benchmark
To assess the model’s generalization capability in multilingual contexts, we selected 38 minor languages with 12 major languages, totaling 50 languages.
  • Major Languages English, Chinese, German, Portuguese, Spanish, French, Russian, Hindi, Italian, Korean, Japanese, Arabic
  • Minor Languages Mongolian, Thai, Vietnamese, Lao, Malagasy, Cebuano, Sundanese, Ilokano, Dogue, Croatian, Galician, Czech, Corsican, Luxembourgish, Latin, Ukrainian, Bosnian, Bulgarian, Esperanto, Maithili, Serbian, Albanian, Slovak, Danish, Sanskrit, Norwegian, Guarani, Scottish Gaelic, Kurdish (Sorani), Maltese, Hebrew, Lingala, Bambara, Swahili, Sepeti, Igbo, Kinyarwanda, Hausa
Minor Languages Dataset
Given the extremely limited data for these minor languages, we extracted 2,000 samples from English data and used Google Translate to create training sets for each language.
Minor Languages Benchmark
For minor languages with limited evaluations, following multilingual evaluation method of Llama3 (Dubey et al., 2024) we translate medical-clinical part of MMLU (Hendrycks et al., 2020) questions and answers using Google Translate. We employed a random selection of 3-shot queries to pose questions to the model, followed by answer extraction and evaluation of the model’s responses.
@@ Appendix B General Experiment Setup
Training
All model use AdamW, a batch size of 32, a sequence length of 4096 and a cosine decay learning rate schedule with a linear warmup of proportion 0.3 for training on 8x NVIDIA A800-SXM4-80GB GPUs. For model initialization and data sampling, we set the random seed to 42. For MoE models, We fine-tune them with the same settings as before after sparse upcycling from a dense model. For all routing strategy in MoE models, router parameters are initialized randomly with a zero-mean normal distribution with standard deviation 0.02. We set the learning rate of the dense model to 1e-4 and the learning rate of the MoE model to 1e-5.
Evaluation
To measure model performance, we extract the options from model output and calculate the accuracy with the reference answer. All evaluations use 3-shot examples. The optimal value is selected based on the average accuracy across three tests for each benchmark.
@@ Appendix C Prompts
C.1 Prompts for Generating QA Pairs from Texts
@@
C.2 Prompt for Generating Doctor-Patient Dialogues
Figure 8: Prompt Template for Generating Doctor-Patient Dialogues
@@ Appendix D Token Expert Routing Construction
Appendix E Different Languages for Hybrid Routing
@@
Appendix F Complete Results on 12 Major and 38 Minor Languages
To comprehensively evaluate the performance of the proposed model across different languages, Tab. 9 presents the accuracy results on 12 major languages and 38 low-resource languages. We showcase the performance of MoE models with various routings (in Sec.3.1.2) and the Apollo-MoE series across different base model sizes (e.g. 0.5B, 1.5B, and 7B parameters in Sec.4) on these languages. The models achieve accuracy on par with that of the major languages, even for low-resource languages, highlighting the model’s strong generalization capabilities when handling resource-scarce languages.
Table 9: Vital Model Perfrmances (Accuracy) on Major and Minor Languages.

### v2

@@ 5 Related Work
Mixture-of-Experts and Sparse Upcycling  Mixture-of-Experts (MoE) models differ from traditional dense models by activating only a subset of parameters, or ”experts”, for each input token. This selective activation reduces computational costs while preserving model capacity. Introduced by Shazeer et al. (2017), MoEs have evolved in models like GShard (Lepikhin et al., 2020), Switch Transformers (Fedus et al., 2022), and Mixtral (Jiang et al., 2024), showing improved performance in large-scale tasks. Recent advancements like Sparse Upcycling (Komatsuzaki et al., 2022) efficiently initialize MoEs by replicating dense model parameters. BTX (Sukhbaatar et al., 2024) further refines MoE efficiency by upcycling specialized models.
Multilingual Medical LLMs  Recent research has focused on expanding multilingual medical data and creating models for linguistically diverse populations. A LLM integrated with machine translation has been proposed to address the scarcity of such datasets (Gangavarapu, 2024). Furthermore, multilingual medical corpora and benchmarks across six languages have been developed (Qiu et al., 2024; Wang et al., 2024). Our work builds on these efforts by investigating multilingual patterns and expanding it to 50 languages.
Multilingual Capability Enhancement Research on Multilingual Models has primarily aimed to enhance multilingual capabilities and understand their mechanisms. GreenPLM (Zeng et al., 2023) shares the same motivation with our work to efficiently expand the model’s multilingual capabilities. Efforts have improved performance through translation (Liang et al., 2023) and cross-lingual alignment (Salesky et al., 2023). Techniques like cross-lingual transfer (Kim et al., 2017) and continuous training in specific languages have further advanced LLMs (Cui et al., 2023), while training from scratch shows potential (Muennighoff et al., 2023). Recent models (Zhao et al., 2023; Nguyen et al., 2024) also exhibit strong multilingual abilities without explicit language alignment. Tang et al. (2024) and Zhao et al. (2024) employ neuron analysis method (Mu & Andreas, 2020) to investigate the mechanism, but they typically focus on fewer than seven languages.
@@ 6 Conclusion
In conclusion, this work advances the development of multilingual medical LLMs by addressing key challenges in data scarcity and model scalability. We first construct a high-quality medical dataset covering 12 high-resource languages. Subsequently, we propose Hybrid-$k$ routing to explore multilingual training in a modular manner in MoE. Based on the Hybrid-$k$ routing, we propose a circuits-based paradigm for interpreting information flow in multilingual contexts. The circuit analysis reveal the ”Spread Out in the End” mechanism, based on which we introduced the Post-MoE architecture and demonstrate its superior multilingual capabilities. Furthermore, by introducing language family experts, we efficiently extend Post-MoE’s capabilities to 50 languages with limited data, demonstrating scalability without additional parameters.
@@
Future Work  In practical applications, the Mixture of Language Family Experts in Post-MoE enables versatile task adaptation and efficient scalability. Additionally, by treating modalities as distinct languages, the proposed framework facilitates studying multimodal interaction dynamics. Theoretically, while many studies have highlighted the presence of multilingual neurons at the bottom layers of the model, in circuit-based paradigm we further observed that activation parameters at bottom layers are concentrated in the Chinese and English language modules rather than being distributed across various language modules and hypothesize that this may reflect a model strategy that relies on core language experts at the bottom layers for multilingual translation. The observation provides an additional dimension (language experts) for analyzing multilingual mechanisms beyond the depth perspective. It prompts further exploration into the selection of language experts at specific depths and their inter-layer interactions.
Acknowledgement
This work was supported by the Shenzhen Science and Technology Program (JCYJ20220818103001002), Shenzhen Doctoral Startup Funding (RCBS20221008093330065), Tianyuan Fund for Mathematics of National Natural Science Foundation of China (NSFC) (12326608), Shenzhen Key Laboratory of Cross-Modal Cognitive Computing (grant number ZDSYS20230626091302006), and Shenzhen Stability Science Program 2023.
@@ A.2 Dataset Procession
Adhering to the established data processing standards, we segmented sentences into chunks, filtered them for medical relevance, and reformatted them into a QA format with prompt shown in App. C.2. The processing details are as follows:
@@
A.3 Benchmark Across 12 High-Resource Languages
@@ A.4 More experimental analysis on the Dataset
For Baselines, We select Gemma2-2B (Team et al., 2024b), Phi-3-mini-4k (Abdin et al., 2024) and Qwen2-7B (Yang et al., 2024) as base models. We simultaneously selected the closed-source models, Claude3 Opus (Anthropic, 2024), and GPT-4o (Achiam et al., 2023), as well as highly competitive open-source models MMed-Llama-3-8B (Qiu et al., 2024), LLama3-OpenBioLLM-8B (Ankit Pal, 2024), and LLama3-8B (Dubey et al., 2024) as baselines. For Training, We use full-parameter fine-tuning with a learning rate of 1e-5 for the 7B and 9B models, and a learning rate of 1e-4 for smaller models. All model use a batch size of 32 and a sequence length of 4096 for training on 8x NVIDIA A800-SXM4-80GB GPUs.
@@
As shown in Tab. 8, the performance of each model significantly improved after training with 12 high-resource languages data, further validating the effectiveness of data quality. The fine-tuned models excelled in the medical domain, significantly surpassing open-source medical models with similar parameter counts. Additionally, the medical model fully fine-tuned based on Gemma2-9B performed nearly as well as large-scale closed-source models.
A.5 Low-Resource Languages Dataset&Benchmark
To assess the model’s generalization capability in multilingual contexts, we selected 38 low-resource languages with 12 high-resource languages, totaling 50 languages.
  • High-Resource Languages English, Chinese, German, Portuguese, Spanish, French, Russian, Hindi, Italian, Korean, Japanese, Arabic
  • Low-Resource Languages Mongolian, Thai, Vietnamese, Lao, Malagasy, Cebuano, Sundanese, Ilokano, Dogue, Croatian, Galician, Czech, Corsican, Luxembourgish, Latin, Ukrainian, Bosnian, Bulgarian, Esperanto, Maithili, Serbian, Albanian, Slovak, Danish, Sanskrit, Norwegian, Guarani, Scottish Gaelic, Kurdish (Sorani), Maltese, Hebrew, Lingala, Bambara, Swahili, Sepeti, Igbo, Kinyarwanda, Hausa
Low-Resource Languages Dataset
Given the extremely limited data for these low-resource languages, we extracted 2,000 samples from English data and used Google Translate to create training sets for each language.
Low-Resource Languages Benchmark
For low-resource languages with limited evaluations, following multilingual evaluation method of Llama3 (Dubey et al., 2024) we translate medical-clinical part of MMLU (Hendrycks et al., 2020) questions and answers using Google Translate. We employed a random selection of 3-shot queries to pose questions to the model, followed by answer extraction and evaluation of the model’s responses.
A.6 Verification of the effectiveness of Google Translate in multilingual medical translation
Table 9: Comparison of GPT-4o and Apollo-MoE Performance on MMLU Medical Sections
Model Ar De En Fr Hi It Pt Zh Avg.
GPT-4o (Medical part of MMLU) 81.4 91.5 91.3 88.3 87.5 87.2 90.2 87.2 88.2
GPT-4o (Translated Medical part of MMLU) 80.9 90.5 91.3 89.6 86.9 87.4 90.8 87.2 88.2
Difference 0.5 1.0 0.0 -1.3 0.6 -0.2 -0.6 0.0 0.0
Apollo-MoE (Medical part of MMLU) 57.9 73.8 74.8 71.5 57.2 72.2 73.4 84.2 68.7
Apollo-MoE (Translated Medical part of MMLU) 58.3 73.5 74.8 72.4 56.9 71.9 73.8 84.5 68.8
Difference -0.4 0.3 0.0 -0.9 0.3 0.3 -0.4 -0.3 -0.1
To verify the accuracy of the translated datasets as suggested, we sampled 100 items from the evaluation sets of two rare languages, Croatian (Hr) and Danish (Da). With the assistance of two Chinese translation professionals in the respective languages, we enlisted three licensed Chinese doctors to annotate the accuracy of the items. The average accuracy rates were 90% and 92%, respectively, partially validating the quality of the translated datasets. Additionally, a prior study (Taira et al., 2021) demonstrated that Google Translate achieves an average accuracy of 82.5% in the complex and rigorous context of emergency medicine.
Meanwhile, for the languages evaluated using MMLU medical parts for evaluation and within MMMLU, we added the accuracy of GPT-4o and Apollo-MoE (develop from Qwen2-7B) on the MMMLU medical part (proofread by OpenAI) and the MMLU translated medical part, and compared the results. As shown in the Tab.9, the models achieved nearly identical scores on these two datasets, with the largest accuracy difference being only 1.3%.
@@ Appendix B General Experiment Setup
Training All model use AdamW, a batch size of 32, a sequence length of 4096 and a cosine decay learning rate schedule with a linear warmup of proportion 0.3 for training on 8x NVIDIA A800-SXM4-80GB GPUs. For model initialization and data sampling, we set the random seed to 42. For MoE models, We fine-tune them with the same settings as before after sparse upcycling from a dense model. For all routing strategy in MoE models, router parameters are initialized randomly with a zero-mean normal distribution with standard deviation 0.02. We set the learning rate of the dense model to 1e-4 and the learning rate of the MoE model to 1e-5.
Evaluation To measure model performance, we extract the options from model output and calculate the accuracy with the reference answer. All evaluations use 3-shot examples. The optimal value is selected based on the average accuracy across three tests for each benchmark.
@@ Appendix C Prompts
C.1 Prompt for Generating Doctor-Patient Dialogues
Prompts for generating doctor-patient dialogues is shown in Fig.7.
Figure 7: Prompt Template for Generating Doctor-Patient Dialogues
C.2 Prompts for Generating QA Pairs from Texts
Prompts for generating QA pairs from texts is shown in Fig.8.
@@ Appendix D Token Expert Routing Construction
Appendix E Advantages of Generalizability
The primary method proposed in this paper, PostMoE with language family experts, offers the advantages of generalizability, enabling adding other languages efficiently and effectively.
Experiments To provide a more intuitive understanding, we selected two additional languages not included in the 50 languages: Bengali (Bn) and Amharic (Am), which belong to the Indo-European and Afro-Asiatic language families respectively, to demonstrate the model’s efficient generalization capability. Specifically, we processed 2,000 data samples for Bengali and Amharic and continued fine-tuning the Dense models and ApolloMoE models.
Results As shown in Tab.10, the experimental results demonstrate the clear advantages of the proposed method in adapting to additional languages. For performance improvement in newly added languages, the ApolloMoE model outperforms the Dense model across all scales, with training-related gains also showing advantages in most cases. For preserving the performance of original languages, the ApolloMoE model maintains or even improves performance across nearly all scales, whereas the Dense model generally experiences a decline in performance.
Table 10: Continue Fine-tuning with Bn and Am Data Based on Dense and PostMoE Models.
Models Avg. Accuracy Lang-Spec. Accuracy
High Low Bn Am
Based on Qwen2-0.5B
Dense 39.2 33.2 36.4 31.5
+Bn 38.2 (-1.0) 33.1 (-0.1) 37.5 (+0.9) -
+Am 38.1 (-1.1) 32.9 (-0.3) - 33.9 (+2.4)
ApolloMoE 40.5 34.6 37.8 33.0
+Bn 40.6 (+0.1) 34.6 (-0.0) 39.7 (+1.9) -
+Am 40.5 (-0.0) 34.6 (-0.0) - 36.0 (+3.0)
Based on Qwen2-1.5B
Dense 52.2 43.7 44.0 35.0
+Bn 49.1 (-3.1) 39.2 (-4.5) 50.8 (+6.8) -
+Am 47.7 (-4.5) 39.1 (-4.6) - 36.4 (+1.4)
ApolloMoE 54.8 44.9 50.4 38.6
+Bn 54.0 (-0.8) 44.9 (-0.0) 55.7 (+5.3) -
+Am 54.8 (-0.0) 44.9 (-0.0) - 41.9 (+3.3)
Based on Qwen2-7B
Dense 69.0 56.7 66.3 35.7
+Bn 68.4 (-1.0) 55.7 (-1.0) 68.9 (+2.6) -
+Am 68.3 (-1.1) 55.3 (-1.4) - 38.2 (+2.5)
ApolloMoE 69.9 58.3 67.1 40.5
+Bn 69.6 (-0.3) 58.5 (+0.2) 69.5(+2.4) -
+Am 69.5 (-0.4) 58.3 (-0.0) - 42.5 (+2.5)
Appendix F Mechanistic Interpretability
Appendix G Different Languages for Hybrid Routing
@@
Appendix H Complete Results on 12 High-Resource and 38 Low-Resource Languages
To comprehensively evaluate the performance of the proposed model across different languages, Tab. 11 presents the accuracy results on 12 high-resource languages and 38 low-resource languages. We showcase the performance of MoE models with various routings (in Sec.3.1.2) and the Apollo-MoE series across different base model sizes (e.g. 0.5B, 1.5B, and 7B parameters in Sec.4) on these languages. The models achieve accuracy on par with that of the high-resource languages, even for low-resource languages, highlighting the model’s strong generalization capabilities when handling resource-scarce languages.
Table 11: Vital Model Performances (Accuracy) on High- and Low-Resource Languages.